### v1

# FlashKAT: Understanding and Addressing Performance Bottlenecks in the Kolmogorov-Arnold Transformer

Abstract
The Kolmogorov-Arnold Network (KAN) has been gaining popularity as an alternative to the multi-layer perceptron (MLP) with its increased expressiveness and interpretability. However, the KAN can be orders of magnitude slower due to its increased computational cost and training instability, limiting its applicability to larger-scale tasks. Recently, the Kolmogorov-Arnold Transformer (KAT) has been proposed, which can achieve FLOPs similar to the traditional Transformer with MLPs by leveraging Group-Rational KAN (GR-KAN). Unfortunately, despite the comparable FLOPs, our characterizations reveal that the KAT is still 123x slower in training speeds, indicating that there are other performance bottlenecks beyond FLOPs. In this paper, we conduct a series of experiments to understand the root cause of the slowdown in KAT. We uncover that the slowdown can be isolated to memory stalls and, more specifically, in the backward pass of GR-KAN caused by inefficient gradient accumulation. To address this memory bottleneck, we propose FlashKAT, which builds on our restructured kernel that minimizes gradient accumulation with atomic adds and accesses to slow memory. Evaluations demonstrate that FlashKAT can achieve a training speedup of 86.5x compared with the state-of-the-art KAT, while reducing rounding errors in the coefficient gradients. Our code is available at https://github.com/OSU-STARLAB/FlashKAT.
1 Introduction
The Kolmogorov-Arnold Network (KAN) has emerged as a popular alternative to the multi-layer perceptron (MLP) liu2024kan . Rather than learning a fixed weight on each edge, the KAN instead learns a flexible non-linearity. In doing so, the KAN is capable of increased expressivity and interpretability compared to the MLP. Initial works have primarily demonstrated the impressive capabilities of the KAN in scientific applications or equation modeling tasks li2025u ; coffman2025matrixkan ; liu2024kan ; koenig2024kan ; wang2024kolmogorov . However, even with the superior performance of the KAN for these tasks, it has fallen short on its applicability to larger-scale tasks in NLP and Computer Vision due to substantially increased parameters and computational cost, lack of GPU hardware optimization (e.g., for basis functions), and training instability yu2024kan ; le2024exploring ; VisionKAN2024 . For example, due to these non-linear edges, an edge may require up to 204 FLOPs compared to an MLP, which would require only 2 FLOPs per edge (1 for multiply and 1 for add) yang2024kolmogorov . Therefore, significant research is needed to reduce the orders of magnitude performance gap between KAN and MLP.
Recently, the Kolmogorov-Arnold Transformer (KAT) has managed to overcome some of the obstacles preventing the KAN from being deployed to large-scale tasks yang2024kolmogorov . It achieves this by replacing each MLP in the Transformer with a Group-Rational KAN (GR-KAN), which augments the input to each linear layer with the safe Pade Activation Unit (PAU)molina2019pad in a grouped fashion (more details in Section 2). In doing so, each GR-KAN would learn a weighted summation of learned nonlinear functions. Unlike the traditional KAN, the GR-KAN is more computationally efficient, reducing the number of FLOPs to only slightly more than that of MLP, and is constructed to better map to GPUs. By modifying the Transformer with GR-KAN, the resulting KAT is capable of achieving impressive results on computer vision tasks.
However, even with a comparable number of FLOPs (and similar expected speed) offered by the KAT, surprisingly, the KAT turns out to be significantly slower compared to its MLP counterpart in our characterization. For instance, our experiments reveal that a KAT (using GR-KAN) is 123x slower during training than an identically sized Vision Transformer (ViT, using MLP) dosovitskiy2020image on Imagenet-1K russakovsky2015imagenet . This prompts us to conduct a series of experiments to fully understand the performance bottleneck of KAT. When isolating the source of the problem, we have identified that the slowdown is caused by a memory bottleneck rather than excessive computation. More specifically, the source of the bottleneck stems from an inefficient method of accumulating gradients for the coefficients of the PAU, which uses atomic adds.
To address the memory bottleneck, we propose FlashKAT, a substantially faster KAT that builds on our optimized method of GR-KAN that not only avoids accumulating gradients with atomic adds but also avoids excessive accesses to high-bandwidth memory (HBM). The method restructures data accesses and allows us to leverage shared memory to perform a sum reduction for gradient accumulations. By implementing our method as a GPU kernel, the resulting FlashKAT can achieve a training speedup of 86.5x compared with the state-of-the-art KAT. Furthermore, we demonstrate that our new kernel can reduce rounding errors in the gradient computations of the PAU coefficients by nearly an order of magnitude, thus improving training stability.
The main contributions of this paper are:
  1. 1.
    Conduct a detailed investigation to understand the computational inefficiencies of the KAT, revealing that the primary performance bottleneck arises not from FLOPs, but from excessive memory traffic caused by inefficient gradient accumulation.
  2. 2.
    Propose FlashKAT, which addresses the identified memory bottleneck by leveraging a substantially more efficient gradient accumulation method to speed up training.
  3. 3.
    Perform an extensive evaluation of FlashKAT, demonstrating significantly reduced training times and rounding errors from the proposed method.
2 Background and Related Works
2.1 Kolmogorov-Arnold Network
Recently, the KAN has been popularized as an alternative to the traditional MLP (liu2024kan, ). It is based on the Kolmogorov-Arnold Theorem, which asserts that a multivariate continuous function on a bounded domain $f:[0,1]^{n}\rightarrow\mathbb{R}$ can be rewritten as a finite summation of univariate continuous functions, $\phi_{q,p}:[0,1]\rightarrow\mathbb{R}$ and $\phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$ as expressed in Equation 1.
$f(x_{1},...,x_{n})=\sum^{2n+1}_{q=1}\phi_{q}(\sum^{n}_{p=1}\phi_{q,p}(x_{p}))$ (1)
Although Equation 1 presents a theoretical form of the Kolmogorov-Arnold Theorem, the authors of (liu2024kan, ) demonstrate its effectiveness in a more generalized form by scaling the width and depth arbitrarily for a defined task. Suppose $n_{l}$ is the number of input nodes for layer $l$, where each input is represented by $x_{l,i}$. Each edge connection to layer $l+1$ has a learned activation $\phi_{l,j,i}(\cdot)$, where $i\in[1,n_{l}]$ and $j\in[1,n_{l+1}]$. Then, each output node is represented by Equation 2.
$x_{l+1,j}=\sum^{n_{l}}_{i=1}\phi_{l,j,i}(x_{l,i})$ (2)
In a vectorized format, we can represent Equation 2 with Equation 3.
$\mathbf{x}_{l+1}=\Phi\circ\mathbf{x}_{l}=\begin{bmatrix}\phi_{l,1,1}(\,\cdot\,)&\cdots&\phi_{l,1,n_{l}}(\,\cdot\,)\\ \vdots&\ddots&\vdots\\ \phi_{l,n_{l+1},1}(\,\cdot\,)&\cdots&\phi_{l,n_{l+1},n_{l}}(\,\cdot\,)\end{bmatrix}\mathbf{x}_{l}$ (3)
Thus, if we have a $L$ layer KAN, it can be represented by Equation 4.
$\text{KAN}(\mathbf{x})=(\Phi_{L}\circ\Phi_{L-1}\circ...\circ\Phi_{1})\mathbf{x}$ (4)
The most widely adopted method for representing $\phi_{l,j,i}(\cdot)$ has been with B-splines liu2024kan , and although the B-spline-based KANs demonstrate promising results, they still have a host of shortcomings yang2024kolmogorov ; yu2024kan . These shortcomings consist of (1) an increased parameter and computational cost compared to the MLP alternative (e.g. increased FLOPs as summarized in Table 1), (2) B-splines not being GPU optimizable (e.g., requiring a recursive algorithm), and (3) training instability in larger networks yang2024kolmogorov . Such shortcomings limit the compatibility of the KAN with the Transformer, restricting its application to larger-scale tasks like computer vision vaswani2017attention .
2.2 Kolmogorov-Arnold Transformer
Unlike its predecessors, littered with shortcomings, the KAT succeeds in merging the KAN into a Transformer. It does so by developing a variation of the KAN called GR-KAN, which (1) shares coefficients across edge groups, reducing FLOPs, (2) leverages safe PAU activation functions, eliminating recursion, and (3) applies a variance-preserving initialization, increasing training stability yang2024kolmogorov . Using the GR-KAN, the authors of yang2024kolmogorov construct the KAT by replacing the MLP with the GR-KAN, where each weight in the MLP is augmented with a non-linear, learned activation function (i.e., a non-linear edge in the KAN). Below are more details.
First, rather than learning $d_{in}\cdot d_{out}$ activations, $\phi_{l,j,i}(\cdot)$, like a KAN, the GR-KAN groups inputs and only learns $\lfloor d_{in}/d_{g}\rfloor$ activations, $F(\cdot)$, where $d_{g}$ is the number of activations that share the same set of learned coefficients in a group. Each output $x_{l+1,j}$ is then a weighted summation of $F_{\lfloor i/d_{g}\rfloor}(x_{l,i})$ using learnable weights, $\mathbf{W}\in\mathbb{R}^{d_{out}\times d_{in}}$. The GR-KAN can be expressed in Equation 5.
[…]
Second, each activation in GR-KAN is a PAU molina2019pad , which we will refer to as a group-wise rational, composed of learnable coefficients $a_{i}$ and $b_{j}$ as shown in Equation 6.
$F(x)=\frac{P(x)}{Q(x)}=\frac{a_{0}+a_{1}x+a_{2}x^{2}+...+a_{m}x^{m}}{1+|b_{1}x+b_{2}x^{2}+...+b_{n}x^{n}|}$ (6)
Third, a variance-preserving initialization is applied for stable learning yang2024kolmogorov . The procedure for applying variance preserving weights requires (1) initializing the coefficients of $F(\cdot)$ to mimic a known activation function (ie. Swish, GeLU, Identity) and (2) initializing the weights in $\mathbf{W}$ according to $\mathcal{N}(0,\frac{\alpha}{d_{in}})$ whereby $\alpha=\frac{\mathbb{E}[\mathbf{F(x)}^{2}]}{Var[\mathbf{x}]}$ with the assumption that $\mathbf{x}\sim\mathcal{N}(0,1)$.
Although the KAT is effective at vision tasks and is more computationally efficient than a ViT using the original KAN for training (a result of overcoming shortcomings (1), (2), and (3)), they are still far short in terms of training speed compared to an identical ViT using an MLP. We demonstrate such results in the next Section.
3 Beyond FLOPs: Understanding Performance Bottlenecks of KAT
Insight 1: KAT is significantly slower than ViT during training.
[…]
Figure 1: Comparison of training time (Fwd+Bwd) for ViT and KAT.
Although the KAT presents a promising alternative to the ViT, it still lags behind the ViT in training time. We illustrate this by comparing the combined forward and backward times for KAT-T (5.7M parameters), KAT-S (22.1M parameters), and KAT-B (86.6M parameters) against identically sized ViT-T, ViT-S, and ViT-B models, using the Imagenet-1k dataset for training russakovsky2015imagenet . For the experiment, each KAT group-wise rational contains 8 groups comprising of 6 numerator and 4 denominator coefficients, and all measurements are taken using an H200 GPU. The results, depicted in Figure 1, show the average time of 100 forward and backward passes, following 5 warmup passes. The data reveals that, for these forward/backward passes, KAT-T is 102x slower than ViT-T, KAT-S is 123x slower than ViT-S, and KAT-B is 116x slower than ViT-B.
Insight 2: FLOPs is not the performance bottleneck.
It is natural to assume that such a dropoff in training speed between the ViT and KAT results from the difference in FLOPs, as the KAT has more FLOPs than the ViT. However, by comparing the estimated FLOPs between the GR-KAN and MLP (Table 1), the only component in the ViT and KAT where they differ, we can dismiss such assumptions.
Table 1: Comparison of parameter counts and FLOPs for different layers. “FuncFLOPs” denotes the FLOPs of the non-linear activation. In KAN, $K$ is the spline order and $G$ the number of intervals; in GR-KAN, $m,n$ are the polynomial degrees and $g$ the number of groups.
Name Number of parameters FLOPs
MLP (ViT) $d_{\mathrm{in}}\times d_{\mathrm{out}}+d_{\mathrm{out}}$ $\mathrm{FuncFLOPs}\times d_{\mathrm{out}}+2\times d_{\mathrm{in}}\times d_{\mathrm{out}}$
KAN $d_{\mathrm{in}}\times d_{\mathrm{out}}\,(G+K+3)+d_{\mathrm{out}}$
$\mathrm{FuncFLOPs}\times d_{\mathrm{in}}+d_{\mathrm{in}}\times d_{\mathrm{out}}$
×$\times\bigl{[}9K\times(G+1.5K)+2G-2.5K+3\,\bigr{]}$
GR-KAN (KAT) $d_{\mathrm{in}}\times d_{\mathrm{out}}+d_{\mathrm{out}}+(m+n\times g)$ $(2m+2n+3)\times d_{\mathrm{in}}+2\times d_{\mathrm{in}}\times d_{\mathrm{out}}$
As can be seen, between the FLOPs expressions of the ViT and KAT, the main difference resides in the MLP activation requiring $\text{FuncFLOPs}\times d_{out}$ FLOPs and the GR-KAN requiring $(2m+2n+3)\times d_{in}$ FLOPs, both of which are negligible compared with the $2\times d_{in}\times d_{out}$ term. This suggests that the slightly increased FLOPs in the KAT may not be the source of the slowdown. To further verify this, we conduct an experiment that artificially increases the FLOPs inside the group-wise rational in the KAT by placing an additional loop around the floating point computations. This would examine whether FLOPs bottleneck the performance. We measure performance metrics using Nvidia Nsight Compute [^1] The documentation for Nvidia Nsight Compute: https://docs.nvidia.com/nsight-compute/ on an RTX 4060 Ti for the forward and backward pass. The dimension configuration is: input tensor, $\mathbf{X}\in\mathbb{R}^{1024\times 197\times 768}$, numerator coefficients, $\mathbf{A}\in\mathbb{R}^{8\times 6}$, denominator coefficients, $\mathbf{B}\in\mathbb{R}^{8\times 4}$, and the output gradient tensor, $\mathbf{dO}\in\mathbb{R}^{1024\times 197\times 768}$. We report these results in Table 2.
Table 2: Performance breakdown for forward and backward passes of the group-wise rational for the KAT. “Loops” is the total number of instructions loops; “Cycles” is the cycle count; “Time” is wall-clock time; “FLOPs” is the floating point operation count; “SM Thp.”, “L1 Thp.”, “L2 Thp.”, and “HBM Thp.” report the percentage of the device’s peak arithmetic and peak memory bandwidth.
Forward Pass
Loops FLOPs Cycles Time SM Thp. (%) L1 Thp. (%) L2 Thp. (%) HBM Thp. (%)
1 2.9T 11.3M 4.89 ms 29.14 28.25 18.11 89.40
2 5.9T 11.3M 4.91 ms 30.34 28.46 18.08 89.25
4 11.8T 11.3M 4.89 ms 41.02 29.08 19.78 89.37
8 23.6T 11.3M 4.91 ms 62.20 28.00 18.40 89.34
Backward Pass
Loops FLOPs Cycles Time SM Thp. (%) L1 Thp. (%) L2 Thp. (%) HBM Thp. (%)
1 11.2T 2.4T 1.03 s 1.97 4.38 5.24 1.01
2 22.3T 2.4T 1.03 s 1.97 4.38 5.22 1.01
4 44.6T 2.4T 1.03 s 1.97 4.38 5.22 1.01
8 89.2T 2.4T 1.03 s 1.97 4.38 5.22 1.01
From Table 2, we can immediately see that, even if the FLOPs scale to 8x of the original amount, the overall required "Cycles" or "Time" to complete the computation does not scale in either the forward or backward pass. Such results clearly demonstrate that the forward and backward pass kernels in KAT are not compute-bound, and we must look beyond FLOPs for performance bottlenecks.
Insight 3: Backward pass dominates execution time.
Results in Table 2 also demonstrate that a far greater amount of compute time is dedicated to the backward pass than the forward pass. For instance, the forward pass takes 4.96 ms, whereas the backward pass requires 1.03 s, which is 207.7x as long. Therefore, in order to improve the training time of the KAT, performance issues for the backward pass computation must be identified and addressed. The backward pass can be broken down into three components (1) loading the values to compute the gradients, (2) computing the gradients, and (3) storing the gradients. Since we know the backward pass is not compute-bound from Insight 2, this means that (1) and (2) are to blame for the slowdown, both of which are memory issues.
Insight 4: Memory (and memory stall in particular) is the culprit.
005005005005001,00010001{,}0001 ,  [^0] Mem Stall - Long ScoreboardMem Stall - LG ThrottleMem Stall - MIO ThrottleMem Stall - MembarMem Stall - BarrierMem Stall - DrainMem Stall - Short ScoreboardMem Stall - Tex ThrottleComputing - Selected981.51981.51981.51981.51680.17680.17680.17680.17454.78454.78454.78454.78387.62387.62387.62387.62220.57220.57220.57220.5764.1164.1164.1164.1120.9520.9520.9520.955.715.715.715.712.382.382.382.38Cycles
Figure 2: The warp-state statistics for the KAT group-wise rational backward pass. "Computing - Selected" is where the warp does useful computation; all the others are memory-stall states.
Focusing on the memory side for backward passes, we note that Table 2 reports very low L1, L2, and HBM throughput. At first glance, this is counterintuitive because a memory-bound process would usually mean a saturated memory bandwidth. However, further reasoning about the underlying hardware can provide insights. A GPU consists of a number of streaming multiprocessors (SMs). Typically, 32 threads are grouped in a warp and executed synchronously. Multiple warps can be scheduled for execution on an SM in a context-switching fashion. When one warp is waiting for a memory transfer, the SM can switch to doing computation for a different "ready" warp (i.e., not waiting for memory), thus hiding the memory latency. However, suppose there are not enough ready warps. Then, the warp scheduler cannot entirely hide long latency or high memory contention, like with atomic operations that create sequential memory accesses. If an SM does not make adequate use of the memory hierarchy consisting of registers, shared memory, and global memory (a mapped memory composed of L1 cache, L2 cache, and HBM), the latency for each transfer becomes even more severe. For reference, on an A100, shared memory, L1, L2, and HBM have a latency of 29.0, 37.9, 261.5, and 466.3 clock cycles, respectively luo2024benchmarking . Assuming that the group-wise rational fails to hide latency, it makes sense that we would witness low memory utilization but still have a memory bottleneck. To validate this assumption, we examine the warp state statistics for the group-wise rational backward pass using Nvidia Nsight Compute with the same configuration as before. These statistics provide the average number of cycles each warp stays in a specific state for each instruction the kernel executes. We report the warp state statistics in Figure 2.
Figure 2 demonstrates that each warp spends a majority of its time waiting for various memory stalls to conclude. Each stall reported is related to a memory transfer, excluding the "Selected" state, where the warp is selected to issue an instruction (i.e., computing). For instance, the "Stall Long Scoreboard", a stall on a memory transfer from global memory, accounts for 412x more time than the time spent in the "Selected" state. As such, it becomes clear that the group-wise rational slowdown is primarily due to memory stalls from memory transfers. This calls for a new method that can be aware of the memory hierarchy and avoid memory operations that create contention and restrict the latency hiding capabilities of the warp scheduler.
4 FlashKAT
In the previous section, we identified that the KAT has a memory bottleneck in the backward pass of the group-wise rational. In this section, we first outline the computations required for this operation, then provide an analysis of the standard group-wise rational used in the state-of-the-art KAT to pin down the source of the bottleneck, and finally present FlashKAT, our proposed Kolmogorov-Arnold Transformer that builds on an optimized group-wise rational that efficiently leverages gradient accumulation and memory hierarchy to minimize global memory accesses.
4.1 Gradient Computations
Since group-wise rational, expressed in Equation 6 as $F(\cdot)$, is an element-wise operation composed of learnable parameters $a_{i}$ and $b_{j}$ for the numerator and the denominator, the backward pass requires us to compute the gradients $\frac{\partial F}{\partial a_{i}}$, $\frac{\partial F}{\partial b_{j}}$, and $\frac{\partial F}{\partial x}$. Suppose that $A(x)=b_{1}x+...,+b_{n}x^{n}$, $\frac{\partial P(x)}{\partial x}=a_{1}+2a_{2}x+...+ma_{m}x^{m-1}$, and $\frac{\partial Q(x)}{\partial x}=\frac{\partial(1+|A(x)|)}{\partial A(x)}\cdot\frac{\partial A(x)}{\partial x}=\frac{A(x)}{|A(x)|}(b_{1}+2b_{2}x+...+nb_{n}x^{n-1})$ then the gradients for each is provided in Equation 7:
$\frac{\partial F}{\partial a_{i}}=\frac{x^{i}}{Q(x)},\quad\frac{\partial F}{\partial b_{j}}=-x^{j}\frac{A(x)}{|A(x)|}\frac{P(x)}{Q(x)^{2}},\quad\frac{\partial F}{\partial x}=\frac{\partial P(x)}{\partial x}\frac{1}{Q(x)}-\frac{\partial Q(x)}{\partial x}\frac{P(x)}{Q^{2}(x)}$ (7)
Equation 7 provides the element-wise gradients; however, since each group of size $d_{g}$, each batch of size $B$ and each element in the sequence of size $N$ share an identical set of coefficients, the contributions of each must be accumulated on $\frac{\partial F}{\partial a_{i}}$ and $\frac{\partial F}{\partial b_{j}}$. Let us denote $\frac{\partial F}{\partial a_{g,i}}$ and $\frac{\partial F}{\partial b_{g,j}}$ as the gradient for coefficient $i$ and $j$ of group $g$. Then we can express this accumulation in Equation 8:
$\frac{\partial F}{\partial a_{g,i}}=\sum^{B}_{i^{\prime}=1}\sum^{N}_{j^{\prime}=1}\sum^{d_{g}}_{k^{\prime}=1}\frac{\partial F(X_{i^{\prime},j^{\prime},g\cdot d_{g}+k^{\prime}})}{\partial a_{g,i}}\text{, and }\frac{\partial F}{\partial b_{g,j}}=\sum^{B}_{i^{\prime}=1}\sum^{N}_{j^{\prime}=1}\sum^{d_{g}}_{k^{\prime}=1}\frac{\partial F(X_{i^{\prime},j^{\prime},g\cdot d_{g}+k^{\prime}})}{\partial b_{g.j}}$ (8)
There are multiple different methods for accumulating these gradients in a kernel, and identifying the optimal accumulation procedure is crucial for a fast training algorithm. We will now demonstrate where the standard group-wise rational method falls short in this regard.
4.2 Analysis of KAT group-wise rational
Provided an input $\mathbf{X}\in\mathbb{R}^{B\times N\times d}$, an output gradient $\mathbf{dO}\in\mathbb{R}^{B\times N\times d}$ and coefficients $\mathbf{A}\in\mathbb{R}^{n_{g}\times m}$ and $\mathbf{B}\in\mathbb{R}^{n_{g}\times n}$, we want to compute the element-wise gradients $\frac{\partial F}{\partial\mathbf{A}}$ ($\mathbf{dA}$), $\frac{\partial F}{\partial\mathbf{B}}$ ($\mathbf{dB})$, and $\frac{\partial F}{\partial\mathbf{X}}$ ($\mathbf{dX}$). In Algorithm 1, we provide the standard method for computing each of these gradients in the KAT. The algorithm is structured sequentially, and all the on-chip computation is done in the subroutine calls, following Equation 7, for readability.
Algorithm 1 The KAT group-wise rational backward pass.
1:$\mathbf{X},\mathbf{dO}\in\mathbb{R}^{B\times N\times d}$,  $\mathbf{A}\in\mathbb{R}^{n_{g}\times m}$,   $\mathbf{B}\in\mathbb{R}^{n_{g}\times n}$ in global memory.
2:Set block size to $S_{\text{block}}$.
3:Divide $\mathbf{X}$ and dO into $T=\lceil(B\cdot N\cdot d/S_{block}\rceil$ blocks $\mathbf{X_{1}},...,\mathbf{X_{T}}$ and $\mathbf{dO_{1}},...,\mathbf{dO_{T}}$ each of size $S_{\text{block}}$.
4:Initialize $\textbf{dA}\leftarrow 0\in\mathbb{R}^{n_{g}\times m}$, $\mathbf{dB}\leftarrow 0\in\mathbb{R}^{n_{g}\times n}$, and $\textbf{dX}\leftarrow 0\in\mathbb{R}^{B\times N\times d}$.
5:for $1\leq i\leq T$ do
6:    Load $\mathbf{X_{i}}$ from global memory
7:    Load $\mathbf{dO_{i}}$ from global memory
8:    for $1\leq j\leq S_{\text{block}}$ do
9:        $k\leftarrow\lfloor((i\cdot T+j)\mod d)/d_{g}\rfloor$
10:        Load $A_{k\cdot m,1},...,A_{k\cdot m,m}$ and $B_{k\cdot n,1},..,B_{k\cdot n,n}$ from global memory
11:        $dX_{i,j}\leftarrow\textbf{Compute\_dX}(A_{k\cdot m,1},...,A_{k\cdot m,m},B_{k\cdot n,1},..,B_{k\cdot n,n},X_{i,j},dO_{i,j})$
12:        $\mathbf{dA_{k\cdot m}}^{\prime}\leftarrow\textbf{Compute\_dA}(A_{k\cdot m,1},...,A_{k\cdot m,m},B_{k\cdot n,1},..,B_{k\cdot n,n},X_{i,j},dO_{i,j})$
13:        $\mathbf{dB_{k\cdot n}}^{\prime}\leftarrow\textbf{Compute\_dB}(A_{k\cdot m,1},...,A_{k\cdot m,m},B_{k\cdot n,1},..,B_{k\cdot n,n},X_{i,j},dO_{i,j})$
14:        Load $\mathbf{dA_{k\cdot m}}$ from global memory, set $\mathbf{dA_{k\cdot m}}\leftarrow\mathbf{dA_{k\cdot m}}+\mathbf{dA^{\prime}_{k\cdot m}}$, write back to global memory
15:        Load $\mathbf{dB_{k\cdot n}}$ from global memory, set $\mathbf{dB_{k\cdot n}}\leftarrow\mathbf{dB_{k\cdot n}}+\mathbf{dB^{\prime}_{k\cdot n}}$, write back to global memory
16:    end for
17:    Write back $\mathbf{dX_{i}}$ to global memory
18:end for
19:return $\mathbf{dA},\mathbf{dB},\mathbf{dX}$
At a high level, Algorithm 1 will iterate over the $T$ blocks in the grid, and for each block there will be $S_{block}$ threads that will compute the associated gradients for $\mathbf{dA}$, $\mathbf{dB}$, $\mathbf{dX}$. In the case of $\mathbf{dA}$ and $\mathbf{dB}$, for each of these computations, the thread will perform an atomic add, which requires reading from global memory, accumulating the value, and then storing the result back to global memory as an indivisible unit of work. However, since each of these threads needs to perform an atomic add for each of the $m+n$ coefficients, the atomic adds must occur sequentially. As such, for each of these atomic adds, a stall occurs. Since multiple threads within the same block and across different blocks may be writing to the same location, there will be extreme resource contention. Such resource contention inevitably leads to high amounts of latency and stalls without utilizing the available bandwidth, as we demonstrated in Section 1.
Let us analyze the total number of global memory accesses incurred by Algorithm 1. We can see that each $X_{i}$ and $dO_{i}$ requires $S_{block}$ reads from global memory. Then writing $dX_{i}$ back to global memory will require $S_{block}$ writes to global memory. Since these 3 memory actions occur $T$ times they will require $3TS_{\text{block}}=3\frac{B\cdot N\cdot d}{S_{\text{block}}}\cdot S_{\text{block}}=3\cdot B\cdot N\cdot d$ global memory accesses. Regarding the coefficients, each element in the output $dX$ requires loading $m+n$ coefficients from global memory. It also requires the aforementioned atomic add, which includes an additional $(m+n)$ reads and $(m+n)$ writes to global memory. Thus, since there are $T$ blocks of $S_{\text{block}}$ threads, there will be $TS_{\text{block}}\cdot 3(m+n)=3(m+n)\cdot B\cdot N\cdot d$ accesses. As a result, there are $3(m+n+1)\cdot B\cdot N\cdot d$ global memory accesses in total. To re-emphasize, the memory accesses for the coefficients for the atomic add are much worse due to the additional contention with other threads.
4.3 FlashKAT group-wise rational
We now present an optimized method for computing the backward pass of the group-wise rational in Algorithm 2. To reduce global memory accesses, we restructure the grid from 1D to 2D. Such a restructuring mirrors the format of Equation 8, allowing an easier mapping of it to the algorithm. The first dimension of our reconstructed grid is allocated for resolving the $B$ and $N$ summations, whereas the second dimension is responsible for the $d_{g}$ summation. The size of this first dimension in the grid is $B\cdot N/S_{\text{block}}$, whereas the size of the second dimension is $n_{g}$ (the total number of groups in GR-KAN).
Algorithm 2 The FlashKAT group-wise rational backward pass.
1:$\mathbf{X},\mathbf{dO}\in\mathbb{R}^{B\times N\times d}$,  $\mathbf{A}\in\mathbb{R}^{n_{g}\times m}$,   $\mathbf{B}\in\mathbb{R}^{n_{g}\times n}$ in global memory.
2:Set block sizes to $S_{block}$ and $d_{g}=d/n_{g}$.
3:Divide $\mathbf{X}$ and $\mathbf{dO}$ into $\lceil(B\cdot N\cdot d/(S_{block}\cdot d_{g})\rceil$ blocks $\mathbf{X_{i,j}}\in\mathbb{R}^{S_{\text{block}}\times d_{g}}$ and $\mathbf{dO_{i,j}}\in\mathbb{R}^{S_{\text{block}}\times d_{g}}$ of a $T\times n_{g}$ grid where $T=\lceil(B\cdot N/S_{block}\rceil$.
4:Divide $\mathbf{A}$ and $\mathbf{B}$ into $n_{g}$ blocks $\mathbf{A_{j}}\in\mathbb{R}^{m}$ and $\mathbf{B_{j}}\in\mathbb{R}^{n}$.
5:Initialize $\mathbf{dA}\leftarrow 0\in\mathbb{R}^{n_{g}\times m}$, $\mathbf{dB}\leftarrow 0\in\mathbb{R}^{n_{g}\times n}$, and $\mathbf{dX}\leftarrow 0\in\mathbb{R}^{B\times N\times d}$.
6:for $1\leq i\leq T$ do
7:    for $1\leq j\leq n_{g}$ do
8:        Load $\mathbf{A_{j}}$, $\mathbf{B_{j}}$ from global memory
9:        Load $\mathbf{X_{i.j}}$ from global memory
10:        Load $\mathbf{dO_{i,j}}$ from global memory
11:        Initialize $\mathbf{dA_{j}}^{\prime}\leftarrow 0$ and $\mathbf{dB_{j}}^{\prime}\leftarrow 0$
12:        for $1\leq k\leq S_{\text{block}}$ do
13:           for $1\leq l\leq d_{g}$ do
14:               $dX_{i,j,k,l}\leftarrow\textbf{Compute\_dX}(A_{j,1},...,A_{j,m},B_{j,1},..,B_{j,n},X_{i,j,k,l},dO_{i,j,k,l})$
15:               $\mathbf{dA_{j}}^{\prime}\leftarrow\mathbf{dA_{j}}^{\prime}+\textbf{Compute\_dA}(A_{j,1},...,A_{j,m},B_{j,1},..,B_{j,n},X_{i,k},dO_{i,j,k,l})$
16:               $\mathbf{dB_{j}}^{\prime}\leftarrow\mathbf{dB_{j}}^{\prime}+\textbf{Compute\_dB}(A_{j,1},...,A_{j,m},B_{j,1},..,B_{j,n},X_{i,j,k,l},dO_{i,j,k,l})$
17:           end for
18:        end for
19:        Load $\mathbf{dA_{j}}$ from global memory, set $\mathbf{dA_{j}}\leftarrow\mathbf{dA_{j}}+\mathbf{dA_{j}}^{\prime}$, write back to global memory
20:        Load $\mathbf{dB_{j}}$ from global memory, set $\mathbf{dB_{j}}\leftarrow\mathbf{dB_{j}}+\mathbf{dB_{j}}^{\prime}$, write back to global memory
21:        Write back $\mathbf{dX_{i,j}}$ to global memory
22:    end for
23:end for
24:return $\mathbf{dA},\mathbf{dB},\mathbf{dX}$
Algorithm 2 does not change the memory accesses for reads and writes for $\mathbf{dX},\mathbf{X}$ and $\mathbf{dO}$; however, it heavily influences them for $\mathbf{A}$, $\mathbf{B}$, $\mathbf{dA}$, and $\mathbf{dB}$. At a high level, we can see that our algorithm iterates over the grid of $T\cdot n_{g}$ blocks. For each of these blocks, we load a single group of coefficients for $\mathbf{A_{j}}$ and $\mathbf{B_{j}}$. Then, we use this single set of coefficients for all computations in the block. As a result, rather than each thread performing an atomic add for each value in $\mathbf{dA_{j}}$ and $\mathbf{dB_{j}}$, the algorithm can instead accumulate all the contributions for $\mathbf{dA_{j}}$ and $\mathbf{dB_{j}}$ in the block before doing an atomic add. As a result, the total amount of memory contention and global memory accesses reduce.
We will now analyze the total global memory accesses for Algorithm 2. A block will first load one set of coefficients from the global memory associated with the group to which the block is assigned. Since there is only 1 set of coefficients per block, this action will only require $m+n$ accesses. Then the algorithm will load the associated $\mathbf{X_{i,j}}$ and $\mathbf{dO_{i,j}}$ from global memory, each of which is of size $S_{\text{block}}\cdot d_{g}$. After computing $\mathbf{dX_{i,j}}$, its $S_{\text{block}}\cdot d_{g}$ values will be written back to global memory. Then, an atomic add will load $m+n$ values from $\mathbf{dA_{j}}$ and $\mathbf{dB_{j}}$, accumulate them with $\mathbf{dA_{j}}^{\prime}$ and $\mathbf{dB_{j}}^{\prime}$, and write them back to global memory. Since all these operations occur across the $T\cdot n_{g}$ blocks, there will be $T\cdot n_{g}(3S_{\text{block}}\cdot d_{g}+3(m+n))=3\frac{B\cdot N}{S_{\text{block}}}\cdot\frac{d}{d_{g}}(S_{\text{block}}\cdot d_{g}+(m+n))=3(\frac{m+n}{S_{\text{block}}\cdot d_{g}}+1)\cdot B\cdot N\cdot d$ global memory access. Compared to Algorithm 1, we can see that Algorithm 2 reduces the number of global memory accesses and subsequently the number of atomic adds by a factor of $\frac{1}{S_{\text{block}}\cdot d_{g}}$.
5 Results
5.1 Revisiting Speed, Utilization and Stalls
Using our experimental configurations from Section 1, we verify the efficacy of Algorithm 2 compared to Algorithm 1. We implement Algorithm 2 in Triton 10.1145/3315508.3329973 and run our experiments on a 4060 Ti. First, we compare the backward pass of the FlashKAT group-wise rational against the KAT group-wise rational by measuring cycles, time, and throughput utilization for SMs and memory. As shown in Table 3, the FlashKAT group-wise rational achieves a speedup of 140.5x over the KAT group-wise rational, with improvement across the utilization of SM, cache, and memory, contributing to the reduction in backward pass time.
Table 3: A comparison between the group-wise rational backward pass for KAT and FlashKAT using FLOPs, cycles, time, SM, and memory throughput.
Model Cycles Time SM Thp. (%) L1 Thp. (%) L2 Thp. (%) HBM Thp. (%)
KAT 2.4T 1.03 s 1.97 4.38 5.24 1.01
FlashKAT 16.9M 7.33 ms 32.24 34.14 44.76 92.05
0011112222Mem Stall - Long ScoreboardComputing - Selected2.312.312.312.311111Cycles
Figure 3: The warp-state statistics for the FlashKAT group-wise rational backward pass. "Computing - Selected" is where the warp does useful computation; all the others are memory-stall states.
Secondly, we conduct again the analysis of warp state statistics for the backward pass of FlashKAT group-wise rational. From observing the warp state statistics in Figure 3, we can see that only the "Stall Long Scoreboard" is still longer than the "Selected" compute state; all the other 8 memory stalls are shorter than compute. Furthermore, the "Stall Long Scoreboard" has reduced from 981.51 cycles to 2.31 cycles. Such a result indicates that the new kernel is now spending more time performing computation rather than mostly waiting for memory transfers.
5.2 Image Recognition Training Speed and Accuracy
We validate that the training capabilities of FlashKAT are consistent with KAT by training it on ImageNet-1k for image recognition (russakovsky2015imagenet, ). While training we also measure the training throughput of FlashKAT to compare against KAT, ViT, and DeiT. The data loader time is excluded from this measurement as it is dependent on external factors (e.g., CPU, main memory, etc.). We measure the throughput in images/second as the average of 100 samples, with 95% confidence intervals provided.
Our training procedure follows an identical curriculum to yang2024kolmogorov , wherein we follow the hyperparameters of DeiT touvron2021training . As such, we train using a batch size of 1024 with the AdamW optimizer loshchilov2017decoupled . In GR-KAN, each of the 8 groups shares its 6 numerator coefficients and has 4 unique denominator coefficients. The attention layers are initialized with Mimetic initialization trockman2023mimetic . The first layer of GR-KAN has its group-wise rational initialized to the identity function, and the second layer is initialized to a Swish function ramachandran2017searching . The patch size is set to 16. For data augmentation and regularization techniques, we apply RandAugment cubuk2020randaugment , Mixup zhang2017mixup , CutMix yun2019cutmix , Random Erasing zhong2020random , weight decay, Label Smoothing szegedy2016rethinking and Stochastic Depth huang2016deep . All experiments are conducted using an H200 GPU. Appendix A.1 provides the complete set of model configurations and training hyperparameters.
Table 4: Comparative analysis of model performance and computational efficiency on ImageNet-1K.
Model #Param. Top-1 Train Thp. (images/s)
ViT-T 5.7 M 72.7 8954.97 ($\pm$ 17.36)
DeiT-T 5.7 M 72.2 8954.97 ($\pm$ 17.36)
KAT-T 5.7 M 74.6 87.73 ($\pm$0.01)
FlashKAT-T 5.7 M 74.6 6317.90 ($\pm$2.65)
ViT-S 22.1 M 78.8 5311.71 ($\pm$ 6.58)
DeiT-S 22.1 M 79.8 5311.71 ($\pm$ 6.58)
KAT-S 22.1 M 81.2 43.28 ($\pm$ 0.01)
FlashKAT-S 22.1 M 81.4 3741.91 ($\pm$ 0.08)
ViT-B 86.6 M 79.1 2457.15 ($\pm$ 2.08)
DeiT-B 86.6 M 81.8 2457.15 ($\pm$ 2.08)
KAT-B 86.6 M 82.3 21.24 ($\pm$ 0.02)
FlashKAT-B 86.6 M 82.2 1801.75 ($\pm$ 0.683)
We provide the image recognition accuracy of FlashKAT-T, FlashKAT-S, and FlashKAT-B models on the ImageNet-1k dataset in Table 4. Importantly, FlashKAT-T, FlashKAT-S, and FlashKAT-B all match their original counterparts and continue to outperform their ViT and DeiT alternatives in image classification. We also report the respective training throughput in Table 4. Our results demonstrate that the proposed FlashKAT-T, FlashKAT-S, and FlashKAT-B are 72.0x, 86.5x, and 84.5x faster than their KAT counterparts. This significantly reduces the performance gap to ViT, making KAT-based models a viable option for vision tasks.
5.3 Reduced Gradient Rounding Error
Table 5: A comparison of mean absolute error (with 95% confidence intervals) and variances.
KAT
Gradient Mean Absolute Error Variance
$\mathbf{dA}$ $8.84\times 10^{-2}\;(\pm 3.40\times 10^{-3})$ $1.45\times 10^{-2}$
$\mathbf{dB}$ $9.63\times 10^{-2}\;(\pm 9.87\times 10^{-3})$ $8.11\times 10^{-2}$
FlashKAT
Gradient Mean Absolute Error Variance
$\mathbf{dA}$ $8.42\times 10^{-4}\;(\pm 3.28\times 10^{-5})$ $1.35\times 10^{-6}$
$\mathbf{dB}$ $9.81\times 10^{-4}\;(\pm 1.15\times 10^{-4})$ $1.11\times 10^{-5}$
In addition to a significant speedup, the proposed FlashKAT also offers reduced rounding error when computing the gradients for the learned group-wise rational coefficients. This is a byproduct of our method, accumulating gradients in each block through a reduction sum rather than using atomic adds. Table 5 demonstrates that the mean absolute error in gradient outputs in FlashKAT is orders of magnitude smaller than KAT. As such, FlashKAT is less susceptible to rounding errors that may contribute towards unstable gradient updates (e.g. low-precision training). Due to space limitations, more details are provided in Appendix A.2.
6 Conclusion
The successes of the KAN, due to its increased expressivity and interpretability, have made it a popular alternative to the MLP. However, it has yet to achieve widespread application in large-scale NLP and Vision tasks due to its increased parameters, computational cost, lack of GPU hardware optimization, and training instability. The KAT has resolved many of the KAN’s shortcomings with the introduction of GR-KAN, but it still suffers from slow training speeds. In this work, we conduct experimentation to understand the source of the slowdown and identify that the slow training speeds are a byproduct of memory stalls in the backward pass. To address this bottleneck, we propose a novel restructuring of the group-wise rational that allows efficient gradient accumulation by minimizing atomic add usage. The resulting FlashKAT achieves an 86.5x speedup compared to ViT, while reducing gradient rounding errors. The application of our work both enables wider adoption and expedites research on the KAT.
Limitations. A main drawback preventing the use of KAN in the Transformer is the increased computation of the KAN. With GR-KAN, the KAT successfully reduced the FLOPs to nearly match the speed of an MLP on the forward pass, but it is still orders of magnitude slower on the backward pass, making training a computational burden. Our proposed FlashKAT, with a significant speedup, has mostly filled the performance gap, but is still currently about 20% slower than a standard Transformer. Even so, our work is an important step in closing the gap between the KAT and standard Transformer in terms of training speed, motivating future work to narrow the gap even further.
Appendix A Appendix
A.1 Image Recognition Training Hyperparameters
We provide the model architecture for KAT-T, KAT-S, and KAT-B in Table 6.
Table 6: Details of KAT model variants.
Model Layers Hidden size (D) MLP size Heads Params
KAT-T 12 192 768 3 5.7 M
KAT-S 12 384 1536 6 22.1 M
KAT-B 12 768 3072 12 86.6 M
We provide the training hyperparameters for KAT-T, KAT-S, and KAT-B in Table 7.
Table 7: Hyperparameters of KAT on ImageNet-1K classification.
Hyperparameter KAT-T KAT-S KAT-B
Input resolution 224×224 224×224 224×224
Epochs 300 300 300
Batch size 1024 1024 1024
Optimizer AdamW AdamW AdamW
Adam $\epsilon$ $1\times 10^{-8}$ $1\times 10^{-8}$ $1\times 10^{-8}$
Adam $(\beta_{1},\beta_{2})$ (0.9, 0.999) (0.9, 0.999) (0.9, 0.999)
Learning rate $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Learning rate decay Cosine Cosine Cosine
Gradient clipping None None None
Warmup epochs 5 5 5
Weight decay 0.05 0.05 0.05
Rand Augment 9/0.5 9/0.5 9/0.5
Repeated augmentation off off off
CutMix 1.0 1.0 1.0
MixUp 0.8 0.8 0.8
CutMix–MixUp switch prob. 0.5 0.5 0.5
Random erasing prob. 0.25 0.25 0.25
Label smoothing 0.1 0.1 0.1
Peak stochastic depth rate 0.1 0.1 0.4
EMA decay rate 0.9999 0.9999 0.9999
A.2 Reduced Gradient Rounding Error
FlashKAT also reduces rounding error when computing the gradients for the learned group-wise rational coefficients, a byproduct of our implementation accumulating gradients in each block through sum reduction rather than using atomic adds. We validate that FlashKAT reduces gradient rounding errors by constructing an experiment measuring the coefficient gradient rounding errors. Our experiment generates $\mathbf{X}\in\mathbb{R}^{1024\times 197\times 768}$, $\mathbf{dO}\in\mathbb{R}^{1024\times 197\times 768}$, $\mathbf{A}\in\mathbb{R}^{8\times 6}$, and $\mathbf{B}\in\mathbb{R}^{8\times 4}$ whose entries are distributed according to $\mathcal{N}(0,1)$. Then we compute $\mathbf{dA}\in\mathbb{R}^{8\times 6}$ and $\mathbf{dB}\in\mathbb{R}^{8\times 4}$ using the KAT group-wise rational backward pass method in float64 and float32, and the FlashKAT group-wise rational backward pass method in float32. We measure the mean absolute error (MAE) over 100 passes between the float64 gradient output and two float32 outputs. All these experiments were done on an RTX 4060 Ti. We report the mean, 95% confidence interval, and variance in Table 8.
Table 8: A comparison of absolute differences in means (with 95% confidence intervals) and variances.
KAT
Gradient Mean Absolute Error Variance
$\mathbf{dA}$ $8.84\times 10^{-2}\;(\pm 3.40\times 10^{-3})$ $1.45\times 10^{-2}$
$\mathbf{dB}$ $9.63\times 10^{-2}\;(\pm 9.87\times 10^{-3})$ $8.11\times 10^{-2}$
FlashKAT
Gradient Mean Absolute Error Variance
$\mathbf{dA}$ $8.42\times 10^{-4}\;(\pm 3.28\times 10^{-5})$ $1.35\times 10^{-6}$
$\mathbf{dB}$ $9.81\times 10^{-4}\;(\pm 1.15\times 10^{-4})$ $1.11\times 10^{-5}$
From Table 8, we can see that the MAE between the KAT float64 and FlashKAT float32 gradient outputs is nearly 2 orders of magnitude smaller compared to the MAE with the KAT float32 gradient outputs. Such a result verifies that our implementation is less susceptible to rounding errors. Although not extremely visible in training at float32, where rounding errors are already minor, the reduction in rounding errors from FlashKAT could be helpful for low-precision training where gradient updates are more unstable.